# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ Abstract
We detail new force-free simulations to investigate magnetosphere evolution and precursor electromagnetic (EM) signals from binary neutron stars. Our simulations fully follow a representative inspiral motion, capturing the intricate magnetospheric dynamics and their impact on EM outflows. We explore a range of stellar magnetic moment orientations and relative strengths, finding that the magnetospheres and Poynting flux evolution are strongly configuration-dependent. The Poynting flux exhibits pulsations at twice the orbital frequency, $2\Omega$, and is highly anisotropic, following a power-law dependence on orbital frequency. The index ranges from 1 to 6, shaped by the intricate dynamics of the magnetospheres. Furthermore, we present the first computation of: (1) The EM forces acting on the star surfaces, revealing the presence of torques that, for highly-magnetized stars, could influence the orbital dynamics or break the crust. (2) The high-energy emission signals from these systems by adopting the established isolated pulsar theory. Assuming curvature radiation in the radiation reaction limit, we find that photons could reach TeV–PeV energies in the last $\sim$ ms for magnetic field strengths $10^{10}-10^{15}$ G. However, our analysis of single photon magnetic pair production suggests that these photons are unlikely to escape, with the MeV band emerging as a promising observational window for precursor high-energy emission. In this framework, we construct high-energy emission skymaps and light curves, exploring observational implications. Finally, we propose potential precursor radio emission and delayed afterglow echoes from magnetized outflows, which may contribute to late-time re-brightening in short gamma-ray bursts or to orphan afterglows.
@@ 1 Introduction
Binary neutron star (BNS) mergers are some of the most extraordinary events of our universe (e.g., E. Burns, 2020). During the final moments of their inspiral, significant amounts of energy ( $\sim 0.01\,M_{\odot}c^{2}$) are liberated in the form of gravitational waves (GWs). The joint detection GW170817/GRB170817 (B. P. Abbott et al., 2017a; A. Goldstein et al., 2017) established the long-predicted connection between the BNS mergers and some gamma-ray bursts (D. Eichler et al., 1989), with the subsequent ‘kilonova’ proving them as r-process sites of heavy elements.
Although the vast majority of the electromagnetic (EM) emission is produced after the merger and consequently, a large number of studies are conducted on that stage, involving both observational discussions and numerical simulations, less attention has been given to precursor EM emission at earlier stages just prior to the merger (see A. G. Suvorov et al., 2024, for a review).
It is well known that pulsars are broadband emitters from radio up to GeV and TeV $\gamma$-rays (D. A. Smith et al., 2023; H. E. S. S. Collaboration et al., 2023). In rotation-powered pulsars, today’s consensus holds that high-energy emission originates predominantly in the equatorial current sheet extending beyond the light cylinder in force-free (FF) configurations (X.-N. Bai & A. Spitkovsky, 2010; I. Contopoulos & C. Kalapotharakos, 2010; C. Kalapotharakos et al., 2014; B. Cerutti et al., 2016; C. Kalapotharakos et al., 2017, 2018; A. A. Philippov & A. Spitkovsky, 2018; G. Cao & X. Yang, 2019; C. Kalapotharakos et al., 2023; B. Cerutti et al., 2024). This region, where strong electric fields accelerate particles, plays a central role in the production of high-energy photons. Additionally, it appears that the efficiency of electron-positron pair production within the magnetosphere regulates the overall efficiency of high-energy emission, acting as a critical factor in determining the observed luminosity and spectral characteristics. Similar physical processes ought to operate in the BNS magnetospheres, which, just prior to coalescence, attain orbital periods of milliseconds. Considered a system in the far field, the conditions due to orbital motion are not too dissimilar from an isolated pulsar with a significant multipolar field structure.
@@
These precursor signals contain additional information about the nature of the merging objects and complement those derived from GWs and the post-merger signals. These include improved localization, binary parameters, viewing geometry and orientation, and NS equation of state.
@@
There are several gamma-ray bursts (GRBs), likely originating from mergers, that exhibit precursors (E. Troja et al., 2010; S.-Q. Zhong et al., 2019; J.-S. Wang et al., 2020; P. Coppin et al., 2020). However, in the absence of joint GW detections, it remains challenging to assess whether they are truly precursors to the GRB, components of the GRB prompt emission, or are associated with phenomena that occur after the merger but prior to the main GRB. Understanding the beaming patterns of precursors and their evolution up to the merger can bring valuable insight into these events.
@@
In this direction, there are a few studies discussing the precursor signals originating from magnetospheric interactions of BNSs. These studies propose estimates for EM luminosities, or Poynting fluxes using simplified considerations, invoking the unipolar inductor model P. Goldreich & D. Lynden-Bell (1969), assuming the simpler case of binaries with one low and one high field companion (B. M. S. Hansen & M. Lyutikov, 2001; A. L. Piro, 2012; D. Lai, 2012; K. N. Gourgouliatos & D. Lynden-Bell, 2019), or in more general binaries of similar fields (V. M. Lipunov & I. E. Panchenko, 1996; M. V. Medvedev & A. Loeb, 2013; M. Lyutikov, 2019; A. J. Cooper et al., 2023). Nonetheless, the rich complexity of the non-trivial magnetospheric interactions and the plethora of the involved parameters (even for circular inspirals) require numerical simulations.
@@
C. Palenzuela et al. (2013a, b) used resistive magnetohydrodynamics (MHD) simulations to study the magnetospheres of inspiral BNSs and calculated the Poynting flux evolution for three cases of magnetic field configurations with aligned, anti-aligned and unequal magnetic moments, followed by cases with perpendicular magnetic moments (M. Ponce et al., 2014), highlighting the magnetosphere diversity and rich phenomenology of their EM emissions. E. R. Most & A. A. Philippov (2020, 2022) using special-relativistic FF electrodynamics simulations, considered cases with fixed separation between the neutron stars (NSs) and explored magnetic field topologies including the anti-aligned case, introducing also inclined dipoles and quadrupole fields on the stars. Their study focused on energy dissipation within BNS magnetospheres and the generation of precursor flares during the late inspiral. More recently, J. F. Mahlmann & A. M. Beloborodov (2024), with similar simulations, investigated the rise of Kelvin-Helmholtz instability at the interface between two NSs with aligned magnetic moments as well as cases where the second star has an inclined magnetic moment, claiming that dissipation occurs even without flux tubes connecting the two stars. In general, such simulations (see also B. Crinquand et al., 2019; F. Carrasco & M. Shibata, 2020) are essential for capturing qualitative details that cannot be addressed by analytic estimates offering quantitative insights that extend beyond the limitations of simple magnetostatic and vacuum approximations (e.g., A. R. King et al., 1990; K. Ioka & K. Taniguchi, 2000; M. Vasúth et al., 2003; M. Lira et al., 2022; C. Aykroyd et al., 2023; R. Tang et al., 2024).
@@
In this work, we comprehensively survey the parameter space of allowed magnetic field configurations of BNS during their inspirals. Using numerical FF MHD simulations, we model the magnetosphere structure, we calculate the outward-propagated EM flux and its sky distribution for a variety of cases involving different star magnetic field ratios and alignments. In addition to the directly observed EM outflows, the non-uniform Poynting flux distribution indicates the potential for “echoes,” where the outward EM flux interacts with material in the local environment and may be observed post-merger.
@@
We also present the first calculation of the backreaction EM torques experienced by each NS component. We analyze our results to determine the evolution of the Poynting flux and its dependence on the orbital frequency up to the merger, finding non-trivial dependence varying for different set of parameters. Furthermore, invoking observationally-validated theoretical approaches from isolated pulsar studies, we investigate the production and emission of high-energy photons and examine the emission patterns and spectral properties.
@@
This paper structure is as follows: in §2, we describe the numerical methods used in this work, along with parameters and simulation setup. In §3, we present our results, beginning with a detailed description of the magnetosphere structure and Poynting flux dependence. We then compute the forces on the NSs and the resulting torques, discussing their potential impact on orbital dynamics. Additionally, we examine high-energy emissions, deriving their beaming patterns and light curves while also addressing gamma-ray attenuation due to magnetic pair production. Finally, in §4, we conclude with a discussion of the observational implications of our results, while §5 provides a concise summary and outlines pathways for advancing future modeling efforts.
@@ 2.1 Force-Free MHD Approximation
In studies of the magnetospheres of isolated NSs (I. Contopoulos et al., 1999; A. Spitkovsky, 2006; S. S. Komissarov, 2006; A. N. Timokhin, 2006; N. Bucciantini et al., 2006; C. Kalapotharakos & I. Contopoulos, 2009; J. Pétri, 2012; G. Cao et al., 2016; T. Akgün et al., 2016; P. Stefanou et al., 2023; D. Ntotsikas et al., 2024), the FF relativistic ideal MHD framework is typically adopted. This approach is well-justified for the global field structure of rotation-powered pulsars because it assumes locally plasma-filled conditions where the energy density of the EM fields dominates over the particle inertia and thermal pressure and that copious pair production occurs. These conditions are generally satisfied by pulsars which are not very close to the curvature pair production death line. Notably, the FF formalism for isolated pulsars has revealed the existence of a characteristic equatorial current sheet extending beyond the light cylinder, which has been identified as the likely site of gamma-ray emission observed by Fermi-LAT (D. A. Smith et al., 2023). This geometric consistency between the FF solution and gamma-ray observations provides compelling evidence for the validity of the FF framework in describing isolated pulsar magnetospheres.
@@
This framework is not only applicable to isolated pulsars but also extends naturally to the magnetospheres of BNSs. During the inspiral phase, the magnetospheres of the two NSs interact dynamically, yet the fundamental conditions underpinning the FF approximation largely persist.
@@
Moreover, the FF framework effectively captures the behavior of the twisted and dynamic magnetic field structures resulting from the relative orbital motion resulting from the intense gravitationally-driven orbital inspiral. These features are critical for understanding key phenomena such as magnetospheric reconnection, which can power EM precursors and contribute to observable signals such as short GRBs, fast radio transients, and echoes.
@@
Near the merger’s final stages, the tidal distortion of the NSs and the ejection of dense matter from potential crust failure will produce localized regions where particle inertia, radiation, and thermal pressure become significant, necessitating full radiation-MHD treatment to capture the dynamics accurately. Such conditions are likely to arise in the immediate vicinity of the distorted NSs. Nevertheless, in this study, we focus exclusively on the FF regime, as it is expected to prevail on larger scales in lower-density regions of the magnetosphere. By emphasizing the EM field structures under the FF approximation, we aim to isolate the dominant large-scale magnetospheric dynamics that are expected to govern much of the inspiral phase and early merger interactions, providing insights into their role in shaping observable EM signals. For this reason, we also neglect general relativistic distortions to fields or modifications to the 4-current or any dynamical spacetime effects that primarily influence near-surface regions at the final stages of the merger. These approximations result in code that is more efficient in the parameter exploration and characterization we study in this work.
@@ 2.2 Key Configuration Parameters
$\boldsymbol{m}_{i}=m_{i}(\cos{\Phi_{i}}\sin{\alpha_{i}}\hat{\textbf{x}}+\sin{\Phi_{i}}\sin{\alpha_{i}}\hat{\textbf{y}}+\cos{\alpha_{i}}\hat{\textbf{z}})$ (1)
@@
The internal magnetic field along the direction $\boldsymbol{\hat{r}}$ at a radial distance $r$ from its center, is specified as a dipole field in flat spacetime [^2] See J. A. Petterson (1974); I. Wasserman & S. L. Shapiro (1983); A. G. Muslimov & A. I. Tsygan (1986); P. L. Gonthier & A. K. Harding (1994); J. Pétri (2013, 2016) for fields in curved spacetime. More generally, a fluid conducting star would also require a toroidal component for stability (S. K. Lander & D. I. Jones, 2009; T. Akgün et al., 2013; R. Ciolfi & L. Rezzolla, 2013).:
@@
The inclination angles ${\alpha_{1},\alpha_{2}}$ can be parameterized in two ways. One approach allows each inclination angle to range from 00 to $\pi$. Alternatively, the inclination angles can be restricted to the range 00 to $\pi/2$, with $\lambda=\boldsymbol{\Omega\cdot m_{1}/\Omega\cdot m_{2}}$, the ratio between the two magnetic moments, taking either positive or negative values. The latter parameterization arises because any configuration with an inclination angle greater than $\pi/2$ is equivalent to one with an inclination angle in the 00 to $\pi/2$ range but with $\lambda$ flipped in sign and a $180^{\circ}$ rotation of the star in the $\Phi$ direction.
For this study, we chose the second parameterization as it provides a more intuitive classification of cases where the magnetic moment components along the orbital axis are either aligned or anti-aligned. By using the signed magnetic moment ratio, positive when the projections of the moments on the $z$-axis point in the same direction and negative otherwise, we can clearly distinguish between these two classes of configurations. This choice also eliminates redundant cases while preserving all unique magnetospheric structures, allowing for a more efficient exploration of the parameter space.
An isolated pulsar located at $\mathbf{r_{s}}$, with a magnetic field B and spinning at an angular frequency $\boldsymbol{\omega}$ induces an electric field at a point $\mathbf{r}$ given by $\textbf{E}_{s}=-[\boldsymbol{\omega}\times(\textbf{r}-\textbf{r}_{s})]\times\textbf{B}/c$. In the circular binary case, the two stars orbit the system’s barycenter with orbital frequency $\boldsymbol{\Omega}$. As a result, an additional electric field will be induced due to this motion,
@@
Two limiting cases can be considered regarding the spin frequencies of the two stars. In the first case, during the final few rotations of their inspiral, the two stars may become tidally synchronized, resulting in spin frequencies such that $\omega_{1}=\omega_{2}=\Omega$. In the second case, $\Omega$ is significantly higher than the spin frequencies of the individual stars [^3] This will not be true if one or both of the NSs is a recycled millisecond pulsar. In this case, the MSP spin is preferentially parallel to the orbital axes. However, formation channels where one or both members of the BNS is a millisecond pulsar require dynamical interactions in clusters, rather than isolated binary evolution (T. M. Tauris et al., 2017; I. Mandel & F. S. Broekgaarden, 2022) where two low-kick core-collapse supernovae occur. For instance, in the Double Pulsar, the fastest member has a spin period of $22$ ms (A. G. Lyne et al., 2004). , allowing the stars to be approximated as effectively irrotational, with $\omega_{1}=\omega_{2}=0$. This second case is likely for NSs (L. Bildsten & C. Cutler, 1992) while the first is for double white dwarfs. Therefore, the simulations presented in this study adopt the irrotational assumption.

### v2

@@ Abstract
We detail new force-free simulations to investigate magnetosphere evolution and precursor electromagnetic (EM) signals from binary neutron stars. Our simulations fully follow a representative inspiral motion, capturing the intricate magnetospheric dynamics and their impact on EM outflows. We explore a range of stellar magnetic moment orientations and relative strengths, finding that the magnetospheres and Poynting flux evolution are strongly configuration dependent. The Poynting flux exhibits pulsations at twice the orbital frequency, $2\Omega$, and is highly anisotropic, following a power-law dependence on orbital frequency. The index ranges from 1 to 6, shaped by the intricate dynamics of the magnetospheres. Furthermore, we present the first computation of (1) the EM forces acting on the star surfaces, revealing the presence of torques that, for highly magnetized stars, could influence the orbital dynamics or break the crust; (2) the high-energy emission signals from these systems by adopting the established isolated pulsar theory. Assuming curvature radiation in the radiation-reaction limit, we find that photons could reach TeV–PeV energies in the last $\sim\rm{ms}$ for magnetic field strengths $10^{10}-10^{15}$ G. However, our analysis of single photon magnetic pair production suggests that these photons are unlikely to escape, with the MeV band emerging as a promising observational window for precursor high-energy emission. In this framework, we construct high-energy emission skymaps and light curves, exploring observational implications. Finally, we propose potential precursor radio emission and delayed afterglow echoes from magnetized outflows, which may contribute to late-time rebrightening in short gamma-ray bursts or to orphan afterglows.
@@ 1 Introduction
Binary neutron star (BNS) mergers are some of the most extraordinary events of our Universe (e.g., E. Burns, 2020). During the final moments of their inspiral, significant amounts of energy ( $\sim 0.01\,M_{\odot}c^{2}$) are liberated in the form of gravitational waves (GWs). The joint detection GW170817/GRB170817 (B. P. Abbott et al., 2017a; A. Goldstein et al., 2017) established the long-predicted connection between the BNS mergers and some gamma-ray bursts (GRBs; D. Eichler et al., 1989), with the subsequent ‘kilonova’ proving them as r-process sites of heavy elements.
Although the vast majority of the electromagnetic (EM) emission is produced after the merger and, consequently, a large number of studies are conducted on that stage, involving both observational discussions and numerical simulations, less attention has been given to precursor EM emission at earlier stages just prior to the merger (see A. G. Suvorov et al., 2024, for a review).
It is well known that pulsars are broadband emitters from radio up to GeV and TeV $\gamma$-rays (H. E. S. S. Collaboration et al., 2023; D. A. Smith et al., 2023). In rotation-powered pulsars, today’s consensus holds that high-energy emission originates predominantly in the equatorial current sheet extending beyond the light cylinder in force-free (FF) configurations (X.-N. Bai & A. Spitkovsky, 2010; I. Contopoulos & C. Kalapotharakos, 2010; C. Kalapotharakos et al., 2014; B. Cerutti et al., 2016; C. Kalapotharakos et al., 2017, 2018; A. A. Philippov & A. Spitkovsky, 2018; G. Cao & X. Yang, 2019; C. Kalapotharakos et al., 2023; B. Cerutti et al., 2025). This region, where strong electric fields accelerate particles, plays a central role in the production of high-energy photons. Additionally, it appears that the efficiency of electron-positron pair production within the magnetosphere regulates the overall efficiency of high-energy emission, acting as a critical factor in determining the observed luminosity and spectral characteristics. Similar physical processes ought to operate in the BNS magnetospheres, which, just prior to coalescence, attain orbital periods of milliseconds. Considered a system in the far-field, the conditions due to orbital motion are not too dissimilar from an isolated pulsar with a significant multipolar field structure.
@@
These precursor signals contain additional information about the nature of the merging objects and complement those derived from GWs and the postmerger signals. These include improved localization, binary parameters, viewing geometry and orientation, and NS equation of state.
@@
There are several GRBs, likely originating from mergers, that exhibit precursors (E. Troja et al., 2010; S.-Q. Zhong et al., 2019; P. Coppin et al., 2020; J.-S. Wang et al., 2020). However, in the absence of joint GW detections, it remains challenging to assess whether they are truly precursors to the GRB, components of the GRB prompt emission, or associated with phenomena that occur after the merger but prior to the main GRB. Understanding the beaming patterns of precursors and their evolution up to the merger can bring valuable insight into these events.
@@
In this direction, there are a few studies discussing the precursor signals originating from magnetospheric interactions of BNSs. These studies propose estimates for EM luminosities, or Poynting fluxes using simplified considerations, invoking the unipolar inductor model P. Goldreich & D. Lynden-Bell (1969), assuming the simpler case of binaries with one low and one high field companion (B. M. S. Hansen & M. Lyutikov, 2001; D. Lai, 2012; A. L. Piro, 2012; K. N. Gourgouliatos & D. Lynden-Bell, 2019), or in more general binaries of similar fields (V. M. Lipunov & I. E. Panchenko, 1996; M. V. Medvedev & A. Loeb, 2013; M. Lyutikov, 2019; A. J. Cooper et al., 2023). Nonetheless, the rich complexity of the nontrivial magnetospheric interactions and the plethora of the involved parameters (even for circular inspirals) require numerical simulations.
@@
C. Palenzuela et al. (2013a, b) used resistive magnetohydrodynamics (MHD) simulations to study the magnetospheres of inspiral BNSs and calculated the Poynting flux evolution for three cases of magnetic field configurations with aligned, antialigned and unequal magnetic moments, followed by cases with perpendicular magnetic moments (M. Ponce et al., 2014), highlighting the magnetosphere diversity and rich phenomenology of their EM emissions. E. R. Most & A. A. Philippov (2020, 2022) using special-relativistic FF electrodynamics simulations, considered cases with fixed separation between the neutron stars (NSs) and explored magnetic field topologies including the antialigned case, introducing also inclined dipoles and quadrupole fields on the stars. Their study focused on energy dissipation within BNS magnetospheres and the generation of precursor flares during the late inspiral. More recently, J. F. Mahlmann & A. M. Beloborodov (2025), with similar simulations, investigated the rise of Kelvin-Helmholtz instability at the interface between two NSs with aligned magnetic moments as well as cases where the second star has an inclined magnetic moment, claiming that dissipation occurs even without flux tubes connecting the two stars. N. Ortiz et al. (2022) adopted an enclosing surface approximation (a perfectly conducting sphere around the area where the two NSs orbit) to simulate the magnetosphere of synchronized NSs in a binary system at constant separation. Although not a physically realistic scenario, this approach enabled them to numerically calculate accurate fields far away from the system center allowing them to calculate the high-energy skymaps and light curves assuming emissions originate around the separatrix layer and the current sheet. In general, such simulations (see also B. Crinquand et al., 2019; F. Carrasco & M. Shibata, 2020) are essential for capturing qualitative details that cannot be addressed by analytic estimates offering quantitative insights that extend beyond the limitations of simple magnetostatic and vacuum approximations (e.g., A. R. King et al., 1990; K. Ioka & K. Taniguchi, 2000; M. Vasúth et al., 2003; M. Lira et al., 2022; C. Aykroyd et al., 2023; R. Tang et al., 2024).
@@
In this work, we comprehensively survey the parameter space of allowed magnetic field configurations of BNS during their inspirals. Using numerical FF MHD simulations, we model the magnetosphere structure, we calculate the outward-propagated EM flux and its sky distribution for a variety of cases involving different star magnetic field ratios and alignments. In addition to the directly observed EM outflows, the nonuniform Poynting flux distribution indicates the potential for “echoes,” where the outward EM flux interacts with material in the local environment and may be observed postmerger.
@@
We also present the first calculation of the backreaction EM torques experienced by each NS component. We analyze our results to determine the evolution of the Poynting flux and its dependence on the orbital frequency up to the merger, finding nontrivial dependence varying for different set of parameters. Furthermore, invoking observationally validated theoretical approaches from isolated pulsar studies, we investigate the production and emission of high-energy photons and examine the emission patterns and spectral properties.
@@
This paper is structured as follows: In §2, we describe the numerical methods used in this work, along with parameters and simulation setup. In §3, we present our results, beginning with a detailed description of the magnetosphere structure and Poynting flux dependence. We then compute the forces on the NSs and the resulting torques, discussing their potential impact on orbital dynamics. Additionally, we examine high-energy emissions, deriving their beaming patterns and light curves while also addressing gamma-ray attenuation due to magnetic pair production. Finally, in §4, we conclude with a discussion of the observational implications of our results, while §5 provides a concise summary and outlines pathways for advancing future modeling efforts.
@@ 2.1 Force-free MHD Approximation
In studies of the magnetospheres of isolated NSs (I. Contopoulos et al., 1999; N. Bucciantini et al., 2006; S. S. Komissarov, 2006; A. Spitkovsky, 2006; A. N. Timokhin, 2006; C. Kalapotharakos & I. Contopoulos, 2009; J. Pétri, 2012; G. Cao et al., 2016; T. Akgün et al., 2016; P. Stefanou et al., 2023; D. Ntotsikas et al., 2024), the FF relativistic ideal MHD framework is typically adopted. This approach is well justified for the global field structure of rotation-powered pulsars because it assumes locally plasma-filled conditions where the energy density of the EM fields dominates over the particle inertia and thermal pressure and that copious pair production occurs. These conditions are generally satisfied by pulsars, which are not very close to the curvature pair-production death line. Notably, the FF formalism for isolated pulsars has revealed the existence of a characteristic equatorial current sheet extending beyond the light cylinder, which has been identified as the likely site of gamma-ray emission observed by Fermi Large Area Telescope (LAT; D. A. Smith et al., 2023). This geometric consistency between the FF solution and gamma-ray observations provides compelling evidence for the validity of the FF framework in describing isolated pulsar magnetospheres.
@@
This framework not only is applicable to isolated pulsars but also extends naturally to the magnetospheres of BNSs. During the inspiral phase, the magnetospheres of the two NSs interact dynamically, yet the fundamental conditions underpinning the FF approximation largely persist.
@@
Moreover, the FF framework effectively captures the behavior of the twisted and dynamic magnetic field structures resulting from the relative orbital motion resulting from the intense gravitationally driven orbital inspiral. These features are critical for understanding key phenomena such as magnetospheric reconnection, which can power EM precursors and contribute to observable signals such as short GRBs, fast radio transients, and echoes.
@@
Near the merger’s final stages, the tidal distortion of the NSs and the ejection of dense matter from potential crust failure will produce localized regions where particle inertia, radiation, and thermal pressure become significant, necessitating full radiation-MHD treatment to capture the dynamics accurately. Such conditions are likely to arise in the immediate vicinity of the distorted NSs. Nevertheless, in this study, we focus exclusively on the FF regime, as it is expected to prevail on larger scales in lower-density regions of the magnetosphere. By emphasizing the EM field structures under the FF approximation, we aim to isolate the dominant large-scale magnetospheric dynamics that are expected to govern much of the inspiral phase and early merger interactions, providing insights into their role in shaping observable EM signals. For this reason, we also neglect general relativistic distortions to fields or modifications to the four-current or any dynamical spacetime effects that primarily influence near-surface regions at the final stages of the merger. These approximations result in code that is more efficient in the parameter exploration and characterization we study in this work.
@@ 2.2 Key Configuration Parameters
[…]
@@
The internal magnetic field along the direction $\boldsymbol{\hat{r}}$ at a radial distance $r$ from its center is specified as a dipole field in flat spacetime [^2] See (J. A. Petterson, 1974; I. Wasserman & S. L. Shapiro, 1983; A. G. Muslimov & A. I. Tsygan, 1986; P. L. Gonthier & A. K. Harding, 1994; J. Pétri, 2013, 2016) for fields in curved spacetime. More generally, a fluid conducting star would also require a toroidal component for stability (S. K. Lander & D. I. Jones, 2009; T. Akgün et al., 2013; R. Ciolfi & L. Rezzolla, 2013).:
@@
The inclination angles $\alpha_{1},\alpha_{2}$} can be parameterized in two ways. One approach allows each inclination angle to range from 0 to $\pi$. Alternatively, the inclination angles can be restricted to the range 0 to $\pi/2$, with $\lambda=\boldsymbol{\Omega}\cdot\boldsymbol{m}_{1}/\boldsymbol{\Omega}\cdot\boldsymbol{m}_{2}$, the ratio between the two magnetic moments, taking either positive or negative values. The latter parameterization arises because any configuration with an inclination angle greater than $\pi/2$ is equivalent to one with an inclination angle in the 0 to $\pi/2$ range but with $\lambda$ flipped in sign and a $180^{\circ}$ rotation of the star in the $\Phi$ direction.
For this study, we chose the second parameterization as it provides a more intuitive classification of cases where the magnetic moment components along the orbital axis are either aligned or antialigned. By using the signed magnetic moment ratio, positive when the projections of the moments on the $z$-axis point in the same direction and negative otherwise, we can clearly distinguish between these two classes of configurations. This choice also eliminates redundant cases while preserving all unique magnetospheric structures, allowing for a more efficient exploration of the parameter space.
An isolated pulsar located at $\mathbf{r_{s}}$, with a magnetic field B and spinning at an angular frequency $\boldsymbol{\omega}$ induces an electric field at a point $\mathbf{r}$ given by $\textbf{E}_{s}=-[\boldsymbol{\omega}\times(\textbf{r}-\textbf{r}_{s})]\times\textbf{B}/c$. In the circular binary case, the two stars orbit the system’s barycenter with the orbital frequency $\boldsymbol{\Omega}$. As a result, an additional electric field will be induced due to this motion,
@@
Two limiting cases can be considered regarding the spin frequencies of the two stars. In the first case, during the final few rotations of their inspiral, the two stars may become tidally synchronized, resulting in spin frequencies such that $\omega_{1}=\omega_{2}=\Omega$. In the second case, $\Omega$ is significantly higher than the spin frequencies of the individual stars [^3] This will not be true if one or both of the NSs is a recycled millisecond pulsar (MSP). In this case, the MSP spin is preferentially parallel to the orbital axes. However, formation channels where one or both members of the BNS is an MSP require dynamical interactions in clusters, rather than isolated binary evolution (T. M. Tauris et al., 2017; I. Mandel & F. S. Broekgaarden, 2022) where two low-kick core-collapse supernovae occur. For instance, in the double pulsar PSR J0737-3039, the fastest member has a spin period of $22$ ms (A. G. Lyne et al., 2004). , allowing the stars to be approximated as effectively irrotational, with $\omega_{1}=\omega_{2}=0$. This second case is likely for NSs (L. Bildsten & C. Cutler, 1992) while the first is for double white dwarfs. Therefore, the simulations presented in this study adopt the irrotational assumption.